OpenStudy (anonymous):

4. One solution of the equation 6x^2 − 19x + 3 = 0 is 1/6 . What is the other solution? (4 points)

OpenStudy (anonymous):

x1=\[x1=(-b -\sqrt{\Delta})/2a=(19-\sqrt{289})/2*6=1/6\]

OpenStudy (anonymous):

x2=(-b +sqrt{Delta})/2a=(19+sqrt{289})/2*6=3

OpenStudy (anonymous):

Delta=b ^{2}-4a*c=(-19)^{2}-4*6*3=289>0

OpenStudy (anonymous):

so solution is 3

OpenStudy (anonymous):

6x^2 = -19x - 3 6x^2 + 19x + 3 = 0 6x^2 + 18x + x + 3 = 0 6x (x + 3) + 1 (x + 3) = 0 (x + 3) (6x + 1) = 0 x + 3 = 0 -------> x = -3 6x + 1 = 0 -----> x = -1/6